### v3

# Learning Joint Graphical Models with Computational Efficiency, Dynamic Regularization, and Adaptation

Abstract
Multi-sourced datasets are common in studies of variable interactions, for example, individual-level fMRI integration, cross-domain recommendation, etc, where each source induces a related but distinct dependency structure. Joint learning of multiple graphical models (i.e., multiple precision matrices) has emerged as an important tool in analyzing such data. Unlike separate learning, joint learning can leverage shared structural patterns across graphs to yield more accurate results. In this paper, we present an efficient and adaptive method named MIGHT (Multi-task Iterative Graphical Hard Thresholding) to estimate multiple graphs jointly. We reformulate the joint model into a series of multi-task learning problems through a column-by-column manner, and solve these problems using a dynamic regularized algorithm based on iterative hard thresholding. This framework is inherently parallelizable and therefore efficient in computation. Theoretically, we derive the non-asymptotic error bound for the resulting estimator. Furthermore, the proposed algorithm is adaptive to heterogeneous column-wise signal strengths: for nodes with strong signals, our estimator achieves improved error bounds and selection consistency adaptively, and also exhibits asymptotic normality—properties rarely explored in existing joint learning methods. The performance of our method is illustrated through numerical simulations and real data analysis on a cancer gene-expression RNA-seq dataset.
Key words : Graphical model, Joint estimation, Multi-task learning, Nodewise regression, Precision matrix estimation, Selection consistency
1 Introduction
This paper analyzes the joint learning of multiple Gaussian graphical models. For $p$ fixed variables, assume that we collect observations drawn from Gaussian distributions across $K$ different classes or datasets, denoted as $X^{(1)}\in\mathbb{R}^{n_{1}\times p},\cdots,X^{(K)}\in\mathbb{R}^{n_{K}\times p}$. Although their corresponding precision matrices, i.e., the inverse covariance matrices, $\Theta^{(1)},\cdots,\Theta^{(K)}\in\mathbb{R}^{p\times p}$ may exhibit dataset-specific heterogeneity, they often share many common support structures Guo et al. (2011). This structural commonality motivates a joint estimation of these graphical models, rather than separate estimations.
Joint graphical models have significant practical implications in data analysis. For instance, in neuroimaging studies, by collecting $n$ fMRI scans from each of the $K$ participants, we can estimate the functional connectivity network (i.e., graphical model) among the $p$ voxels for each participant. The networks among participants have some shared common structures, indicating functional correspondence. Leveraging these patterns, we can get improved estimation through joint modeling approaches Tsai et al. (2022); van den Bergh et al. (2025). Additionally, transfer learning of graphical models has become a heated topic recently. In this context, joint estimation, which identifies shared structural patterns, can serve as a pre-processing step for transfer learning to assess the similarity between source studies and the target study Li et al. (2023); Zhao et al. (2024). Moreover, joint graphical model estimation plays a significant role in gene expression analysis Lee and Liu (2015), stock price analysis Yang and Peng (2020), graph neural networks Jiang et al. (2022), and so on. These findings necessitate the study of joint graph estimation.
In this paper, we propose an efficient and adaptive method for joint graphical model learning. Through theoretical analysis and practical applications, we demonstrate the advantages of our method in statistical inference.
1.1 Related works
Joint graphical model  Joint graphical model estimation has garnered significant attention over the past decade due to its ability to leverage shared structural patterns across multiple datasets. Pioneering work like Guo et al. (2011) introduced hierarchical regularization to decompose and identify the common shared and class-specific edges among these precision matrices. Danaher et al. (2014) proposed two regularization methods: the Group Graphical Lasso (GGL) and the Fused Graphical Lasso (FGL). GGL combines an element-wise sparsity penalty, $\sum_{i\neq j}\sum_{k\in[K]}|\Theta_{ij}^{(k)}|$, with a covariate-wise penalty, $\sum_{i\neq j}\left\{\sum_{k\in[K]}\left(\Theta_{ij}^{(k)}\right)^{2}\right\}^{1/2}$, to encourage shared sparsity patterns among different graphical models. In contrast, FGL combines the same element-wise sparsity penalty with a fused-type penalty, $\sum_{k<k^{\prime}}\sum_{i\neq j}\left|\Theta_{ij}^{(k)}-\Theta_{ij}^{(k^{\prime})}\right|$, to encourage shared edge values among different graphical models. These two foundational methods have inspired lots of advanced joint estimation methods. Price et al. (2015) proposed a ridge penalty plus a ridge fused penalty to maintain the similarity among the multi-graphs. Gibberd and Nelson (2017) proposed the group-fused graphical lasso method to estimate the time-evolving and piecewise-constant Gaussian graphical model. Bilgrau et al. (2020) extended the work of Price et al. (2015) to the case with prior information. And Navarro et al. (2022) proposed a convex optimization method with $\ell_{1}$ fused penalty under the general case of graph stationarity. Additionally, joint estimation methods based on the CLIME estimator (proposed by Cai et al. (2011)) Lee and Liu (2015); Qiu et al. (2015); Cai et al. (2016), and Bayesian frameworks Li et al. (2019); Jalali et al. (2023), have also been investigated. For a comprehensive review, we refer the reader to Tsai et al. (2022).
Motivation and inspiration   While the aforementioned joint estimation methods capture shared structures, some critical challenges remain: (i) high computational cost, (ii) complex tuning procedures, and (iii) the lack of theoretical guarantees for selection consistency and inference. To address these issues, we begin by considering a group-structured linear model
$y$ $=\sum_{j=1}^{p}X_{G_{j}}\bm{\beta}_{G_{j}}+\epsilon,\qquad\text{with }\underbrace{\sum_{j=1}^{p}\mathbf{1}(\bm{\beta}_{G_{j}}\neq\mathbf{0})\leq s}_{\text{group-wise sparsity}}\text{ and }\underbrace{\sum_{j=1}^{p}\|\bm{\beta}_{G_{j}}\|_{0}\leq ss_{0}}_{\text{element-wise sparsity}},$ (1)
with the covariates partitioned into $p$ groups $\{G_{j}\}_{j=1}^{p}$. Within this joint sparse framework, Simon et al. (2013) introduced the sparse group lasso, which simultaneously applies both an element-wise penalty $\|\bm{\beta}\|_{1}=\sum_{j=1}^{p}\|\bm{\beta}_{G_{j}}\|_{1}$ and a group-wise penalty $\sum_{j=1}^{p}\|\bm{\beta}_{G_{j}}\|_{2}$. The penalty terms of the sparse group lasso and the GGL exhibit similar patterns, reflecting the analogous multiple sparsity structures in their underlying statistical models. Recently, there have been many cutting-edge works on the joint sparse model (1): Cai et al. (2022) provided theoretical guarantees for parameter estimation of the sparse group lasso method. Zhang et al. (2024) developed a double sparse iterative hard thresholding procedure to achieve the minimax adaptivity. Abramovich (2024) proposed a classifier based on logistic regression that leverages such a joint structure (1). Tugnait (2021, 2024) extended the sparse-group-lasso-type penalty to learning multi-attribute graphical models and matrix-valued graphical models, respectively. Motivated by these developments, we introduce an analytical framework that reformulates the joint graphical model into a series of linear models via the nodewise regression approach Meinshausen and Bühlmann (2006); Janková and van de Geer (2017). Then, by developing a dynamic regularized procedure, we can obtain a more refined estimation while maintaining computational efficiency. Our method overcomes the limitations of current joint graphical estimation methods.
1.2 Our contributions
The main contributions of this work are threefold:
  • We propose a computationally efficient method called MIGHT (Multi-task Iterative Graphical Hard Thresholding) to jointly estimate multiple graphs with common structures. We reformulate the joint graphical model estimation problem as a series of multi-task learning problems, and then solve them using a dynamic regularized algorithm based on the iterative hard thresholding. This transformation enables parallel computation, enhancing its efficiency and scalability in practical applications.
  • We establish theoretical guarantees for the MIGHT method. We derive sharper estimation error bounds compared to existing methods and prove that MIGHT achieves selection consistency under proper signal conditions. Additionally, we demonstrate the asymptotic properties of MIGHT, a key feature for statistical inference that is typically absent in regularization-based method literature.
  • We implement the MIGHT method in an open-source R package ADSIHT, available at https://cran.r-project.org/web/packages/ADSIHT/index.html. The comprehensive numerical experiments on synthetic and real datasets demonstrate that our method achieves superior empirical performance than the state-of-the-art methods.
1.3 Organization and Notations
The current paper is organized as follows. Section 2 introduces our method and algorithm. Section 3 derives the theoretical guarantees of our method. Sections 4 and 5 demonstrate the advantages of our method through simulations and empirical data, respectively. Section 6 concludes this paper. All technical proofs and additional simulation results are provided in the appendices and the supplementary material.
Throughout, for the given sequences $a_{n}$ and $b_{n}$, we say that $a_{n}=O(b_{n})$ when $a_{n}\leq Cb_{n}$ for some constant $C>0$, while $a_{n}=o(b_{n})$ corresponds to ${a_{n}}/{b_{n}}\rightarrow 0$ as $n\to\infty$. We write $a_{n}\asymp b_{n}$ if $a_{n}=O(b_{n})$ and $b_{n}=O(a_{n})$. Let $[m]$ denote the set $\{1,2,\cdots,m\}$, and $\mathbf{1}(\cdot)$ as the indicator function. For a set $S$, let $|S|$ denote its cardinality. Let $\mathbf{0}_{p}$ denote the $p$-dimensional zero vector. For a vector $\bm{\beta}\in\mathbb{R}^{p}$, we denote by $\|\bm{\beta}\|_{2}$ its Euclidean norm, by $\text{supp}(\bm{\beta})=\left\{j\in[p]:\beta_{j}\neq 0\right\}$ its support index set, and by $\|\bm{\beta}\|_{0}=|\text{supp}(\bm{\beta})|$ the number of its nonzero entries. For every set $S\subset[p]$, denote by $\bm{\beta}_{S}=(\beta_{j},j\in S)\in\mathbb{R}^{|S|}$ the subvector of $\beta$ indexed by $S$. For a matrix $A$, we denote by $\|A\|_{F}$ its Frobenius norm, and by $\Lambda_{\max}(A)$ and $\Lambda_{\min}(A)$ its largest and smallest eigenvalues. Let $\mathbf{I}_{p}\in\mathbb{R}^{p\times p}$ denote the identity matrix. We use $C,C_{0},C_{1},\cdots$ to denote absolute constants, whose actual values vary from time to time. In the asymptotic analysis, we assume $n,p,K\to\infty$, simultaneously.
2 Methods
This section introduces the model setup and transforms the joint graphical model into a series of multi-task learning problems. A dynamic-regularized iterative hard thresholding-type algorithm is then introduced to solve these problems.
2.1 Model setup
Consider the joint estimation of the graphical models across $K$ independent datasets. For the $k$-th dataset, observations are organized into an $n_{k}\times p$ matrix $X^{(k)}=\left(x_{1}^{(k)},\cdots,x_{n_{k}}^{(k)}\right)^{\top}$, where each row $x_{\ell}^{(k)}$ follows an independent $p$-dimensional Gaussian distribution $N(\bm{\mu}^{(k)},\Sigma^{(k)})$, for every $\ell\in[n_{k}]$. Without loss of generality, we assume $\bm{\mu}^{(k)}=\mathbf{0}_{p}$. Denote by $\Theta^{(k)}:=\left(\Sigma^{(k)}\right)^{-1}\in\mathbb{R}^{p\times p}$ the underlying precision matrix of the $k$-th dataset. For $i\neq j$, the entry $\Theta_{i,j}^{(k)}$ quantifies the conditional dependence between nodes $i$ and $j$ in the $k$-th graph, with $\Theta_{i,j}^{(k)}=0$ indicating conditional independence of $i$ and $j$ given the remaining p2p-2 covariates Lauritzen (1996). We denote by $S_{j}^{(k)}$ the neighbor set of node $j$ in the $k$-th graphical model, i.e.,
$S_{j}^{(k)}:=\left\{i\in[p]\setminus\{j\}:\Theta_{i,j}^{(k)}\neq 0\right\},$
and let $S_{j}:=\bigcup_{k=1}^{K}S_{j}^{(k)}$ denote the total neighbor set of node $j$ across all $K$ graphs, which includes nodes that connect to node $j$ in at least one graph. Define $N:=\sum_{k\in[K]}n_{k}$. We assume that these $K$ graphical models share some common structures, allowing joint estimation to achieve higher accuracy and lower computational cost than separate estimation Guo et al. (2011); Danaher et al. (2014); Tsai et al. (2022). As illustrated in Figure 1, the joint graphs exhibit two structural properties: local edge-sparsity and local neighbor-similarity, which are formalized in the following.
Figure 1: The joint graphical models with $K=4$ graphs and $p=8$ nodes. (a) Edges in these graphs can be divided into three categories: edges universally shared across all graphs (black lines), edges partially shared among subsets (green lines), and edges unique to individual graphs (red lines). (b) Corresponding adjacency matrices for (a). Take node 1 as an example, its neighbor sets across the four graphs are $S_{1}^{(1)}=\{4,5,7\},S_{1}^{(2)}=\{2,4,5\},S_{1}^{(3)}=\{4,5\}$, and $S_{1}^{(4)}=\{5,7\}$. The total neighborhood set $S_{1}=\bigcup_{k=1}^{4}S_{1}^{(k)}=\{2,4,5,7\}$ includes all nodes that connect to node 1 in at least one graph. On average, each edge that is connected to node 11 appears in $\frac{\sum_{k=1}^{4}|S_{1}^{(k)}|}{|S_{1}|}=2.5$ graphs, reflecting the degree of (local) similarity among these graphs.
Definition 1 (Local edge-sparsity)
We say the joint graphical model $\left(\Theta^{(1)},\cdots,\Theta^{(K)}\right)$ is local edge-sparse with parameters $\{s_{j}\}_{j\in[p]}$ if
$|S_{j}|=s_{j},\quad\text{for every }j\in[p].$
We quantify the local edge-sparsity by constraining the total neighbor set for each node $j$ respectively (e.g., $s_{1}=4$ in Figure 1). It also provides a sparse neighbor set in each individual graphical model (since $|S_{j}^{(k)}|\leq|S_{j}|$ for each $(j,k)\in[p]\times[K]$).
Definition 2 (Local neighbor-similarity)
We say the joint graphical model $\left(\Theta^{(1)},\cdots,\Theta^{(K)}\right)$ is local neighbor-similar with parameters $\{s_{0,j}\}_{j\in[p]}$ if
$\frac{\sum_{k=1}^{K}|S_{j}^{(k)}|}{|S_{j}|}=s_{0,j},\quad\text{for every }j\in[p].$
We quantify the local neighbor-similarity by constraining the average frequency of each neighbor for every node $j\in[p]$. An $s_{0,j}$ approaching $K$ allows the near-complete neighbor overlap across all graphs, indicating locally homogeneous structure for node $j$. In contrast, a smaller $s_{0,j}$ encourages each graph to possess some individual neighbors to node $j$, demonstrating greater local heterogeneity among these graphs.
Assume that the joint graphical model exhibits both local edge-sparsity (with $\{s_{j}\}_{j\in[p]}$) and local neighbor-similarity (with $\{s_{0,j}\}_{j\in[p]}$), which we denote as the joint $(s_{j},s_{0,j})_{j\in[p]}$ structure. We next estimate the precision matrices in a column-by-column fashion, i.e., estimating all the $j$-th columns $\left(\Theta_{\cdot,j}^{(1)},\cdots,\Theta_{\cdot,j}^{(K)}\right)$ at once. This is inspired by the nodewise regression Meinshausen and Bühlmann (2006), which reformulates the estimation of a column in a precision matrix as a sparse linear regression problem Janková and van de Geer (2017). For each $(j,k)\in[p]\times[K]$, let $X_{j}^{(k)}\in\mathbb{R}^{n_{k}\times 1}$ denote the observation vector corresponding to the $j$-th covariate, and let $X_{\setminus j}^{(k)}\in\mathbb{R}^{n_{k}\times(p-1)}$ denote the submatrix of $X^{(k)}\in\mathbb{R}^{n_{k}\times p}$ excluding the $j$-th covariate. The next proposition indicates that the column-wise estimation is equivalent to a multi-task learning model.
Proposition 1
The joint estimation of all the j-th column $\left(\Theta_{\cdot,j}^{(1)},\cdots,\Theta_{\cdot,j}^{(K)}\right)$ is equivalent to the multi-task learning model:
$\underbrace{\begin{pmatrix}X_{j}^{(1)}\\ \vdots\\ X_{j}^{(K)}\end{pmatrix}}_{=:X_{j}\in\mathbb{R}^{N}}=$ $\underbrace{\begin{pmatrix}X_{\setminus j}^{(1)}&\cdots&\mathbf{O}_{n_{1}\times(p-1)}\\ \vdots&\ddots&\vdots\\ \mathbf{O}_{n_{K}\times(p-1)}&\cdots&X_{\setminus j}^{(K)}\end{pmatrix}}_{=:X_{\setminus j}\in\mathbb{R}^{N\times(p-1)K}}\underbrace{\begin{pmatrix}\bm{\alpha}_{\setminus j}^{(1)}\\ \vdots\\ \bm{\alpha}_{\setminus j}^{(K)}\end{pmatrix}}_{=:\bm{\alpha}_{\setminus j}\in\mathbb{R}^{(p-1)K}}+\underbrace{\begin{pmatrix}\epsilon_{j}^{(1)}\\ \vdots\\ \epsilon_{j}^{(K)}\end{pmatrix}}_{=:\epsilon_{j}\in\mathbb{R}^{N}},$ (2)
where $N=\sum_{k\in[K]}n_{k}$, and
$\bm{\alpha}_{\setminus j}^{(k)}$ $=-\left(\Theta_{j,j}^{(k)}\right)^{-1}\Theta_{\setminus j,j}^{(k)}\in\mathbb{R}^{p-1},$ (3)
$\epsilon_{j}^{(k)}$ $=X_{j}^{(k)}-X_{\setminus j}^{(k)}\bm{\alpha}_{\setminus j}^{(k)}\sim N\left(\mathbf{0}_{n_{k}},\left(\Theta_{j,j}^{(k)}\right)^{-1}\cdot\mathbf{I}_{n_{k}}\right).$
Since $S_{j}^{(k)}=\text{supp}\left(\Theta_{\setminus j,j}^{(k)}\right)=\text{supp}\left(\bm{\alpha}_{\setminus j}^{(k)}\right)$, after the column-wise transformation (2), the joint $(s_{j},s_{0,j})_{j\in[p]}$ structure is encoded to the coefficient vectors $\left\{\bm{\alpha}_{\setminus j}\right\}_{j\in[p]}$. Additionally, to facilitate subsequent algorithmic analysis, we apply column-wise scaling in model (2), ensuring that each column of $X_{\setminus j}$ has an $L_{2}$-norm of order $\sqrt{N}$. The scaled model is then expressed as
$X_{j}=Z_{\setminus j}\bm{\beta}_{\setminus j}+\epsilon_{j}\in\mathbb{R}^{N},\text{ for each }j\in[p],$ (4)
where $\bm{\beta}_{\setminus j}\in\mathbb{R}^{(p-1)K}$ is composed of $\bm{\beta}_{\setminus j}^{(1)},\cdots,\bm{\beta}_{\setminus j}^{(K)}\in\mathbb{R}^{p-1}$, see Algorithm 2 for specific rescale operations. Therefore, our main objective is to estimate the precision matrices $\left(\Theta^{(1)},\cdots,\Theta^{(K)}\right)$ in a column-by-column fashion, leveraging both local neighbor-similarity and edge-sparsity to achieve accurate and computationally efficient joint estimation.
2.2 The MIGHT method
This subsection introduces a nodewise Multi-task Iterative Graphical Hard Thresholding (MIGHT) to solve the scaled model (4). The key to this method is a two-step thresholding operator $\mathcal{T}_{\lambda,s_{0}}:\mathbb{R}^{(p-1)K}\to\mathbb{R}^{(p-1)K}$ inspired by Zhang et al. (2024), which can be constructed from the following definitions.
  1. 1.
    For a vector $\bm{\beta}=\left(\left(\bm{\beta}^{(1)}\right)^{\top},\cdots,\left(\bm{\beta}^{(K)}\right)^{\top}\right)^{\top}\in\mathbb{R}^{(p-1)K}$, where each subvector $\bm{\beta}^{(k)}\in\mathbb{R}^{p-1}$ corresponds to the $k$-th dataset, we define the edge-wise thresholding operator $\mathcal{T}_{\lambda}^{(1)}:\mathbb{R}^{(p-1)K}\to\mathbb{R}^{(p-1)K}$ as: { T_λ^(1)(β)}_i^(k) = β_i^(k) ⋅1 (|β_i^(k)| ≥λ) , for each $(i,k)\in[p-1]\times[K]$, where $\bm{\beta}_{i}^{(k)}$ represents the $i$-th entry of $\bm{\beta}^{(k)}$.
  2. 2.
    We define the neighbor-wise thresholding operator $\mathcal{T}_{\lambda,s_{0}}^{(2)}:\mathbb{R}^{(p-1)K}\to\mathbb{R}^{(p-1)K}$ as: { T_λ,s_0^(2)(β)}^(k)_i = β^(k)_i ⋅1 { ∑_k’ ∈[K] ( β^(k’)_i )^2 ≥s_0 λ^2 } , for each $(i,k)\in[p-1]\times[K]$.
The two-step hard thresholding can be described as $\mathcal{T}_{\lambda,s_{0}}:=\mathcal{T}_{\lambda,s_{0}}^{(2)}\circ\mathcal{T}_{\lambda}^{(1)}$, i.e., $\mathcal{T}_{\lambda,s_{0}}(\bm{\beta}):=\mathcal{T}_{\lambda,s_{0}}^{(2)}\left(\mathcal{T}_{\lambda}^{(1)}(\bm{\beta})\right)$. Specifically, in the multi-task learning problem (4) with respect to node $j$, for each potential neighbor $i\in[p]\setminus\{j\}$, if $\sum_{k\in[K]}\left\{\left(\hat{\bm{\beta}}_{i}^{(k)}\right)^{2}\cdot\mathbf{1}\left(|\hat{\bm{\beta}}_{i}^{(k)}|\geq\lambda\right)\right\}\geq s_{0}\lambda^{2}$ holds, node $i$ is selected as a neighbor of node $j$, and its edge takes the thresholded value
$\left\{\mathcal{T}_{\lambda,s_{0}}(\hat{\bm{\beta}})\right\}_{i}^{(k)}=\hat{\bm{\beta}}_{i}^{(k)}\cdot\mathbf{1}\left(|\hat{\bm{\beta}}_{i}^{(k)}|\geq\lambda\right),\text{ for each }k\in[K].$
This thresholding mechanism simultaneously controls the edge-sparsity and neighbor-similarity across multiple graphs.
We propose Algorithm 1 to solve the scaled model (4). Starting from the least-squares loss $\mathcal{L}(\bm{\beta}_{\setminus j})=\frac{2}{N}\left\|X_{j}-Z_{\setminus j}\bm{\beta}_{\setminus j}\right\|_{2}^{2}$, in each iteration round the algorithm performs a gradient-descent update and then applies the operator $\mathcal{T}_{\lambda_{t},s_{0}}$ to enforce the desired joint structure. It consists of two stages: the dynamic thresholding iteration and the fixed thresholding iteration. In the dynamic stage (lines 3-7), the thresholding parameter $\lambda_{t}$ decreases geometrically from an initial value $\lambda_{0}$ to a pre-specified fixed value $\lambda_{\infty}$, which effectively balances Type I and Type II errors and yields an initial estimation of $\bm{\beta}_{\setminus j}$. Here, $\kappa\in(0,1)$ is a fixed decreasing rate, typically chosen as 0.9 in practice, and the explicit form of constants $C_{1}^{\text{alg}},C_{2}^{\text{alg}}$ will be given in Theorem 1. In the fixed stage (lines 10-13), we construct an appropriate thresholding parameter $\lambda_{\text{fix}}$, which remains fixed in subsequent iterations. This stage refines the initial estimation, ensuring selection consistency and asymptotic properties.
Algorithm 1 Nodewise Procedure
0:  $Z_{\setminus j},~X_{j},~\kappa,~s_{0,j}$
1:  Initialize $t=0,~\lambda_{\infty}=C_{1}^{\text{alg}}\sqrt{\frac{\log K+(1/s_{0,j})\log p}{N}}$, $\hat{\bm{\beta}}_{\setminus j}^{0}=\mathbf{0}_{(p-1)K}$ and $N=\sum_{k\in[K]}n_{k}$
2:  Initialize $\lambda_{0}=C_{2}^{\text{alg}}$
3:  while $\lambda_{t}\geq\lambda_{\infty},$~ do
4:   $\hat{\bm{\beta}}_{\setminus j}^{t+1}=\mathcal{T}_{\lambda_{t},s_{0}}\left(\hat{\bm{\beta}}_{\setminus j}^{t}+\frac{1}{N}Z_{\setminus j}^{\top}(X_{j}-Z_{\setminus j}\hat{\bm{\beta}}_{\setminus j}^{t})\right)$
5:   $\lambda_{t+1}=\kappa\lambda_{t}$
6:   $t=t+1$
7:  end while
8:  $\hat{s}=\max\left\{1,~\left|\bigcup_{k\in[K]}\text{supp}\left(\left(\hat{\bm{\beta}}_{\setminus j}^{t}\right)^{(k)}\right)\right|\right\}$
9:  $\lambda_{\text{fix}}=C_{1}^{\text{alg}}\sqrt{\frac{\log(\hat{s}K)+(1/s_{0,j})\log p}{N}}$
10:  while $t\leq\left\lceil\frac{\log(\lambda_{0}/\lambda_{\infty})}{\log(1/\kappa)}\right\rceil+C_{4}\log N,$~ do
11:   $\hat{\bm{\beta}}_{\setminus j}^{t+1}=\mathcal{T}_{\lambda_{\text{fix}},s_{0,j}}\left(\hat{\bm{\beta}}_{\setminus j}^{t}+\frac{1}{N}Z_{\setminus j}^{\top}(X_{j}-Z_{\setminus j}\hat{\bm{\beta}}_{\setminus j}^{t})\right)$
12:   $t=t+1$
13:  end while
13:  $\hat{\bm{\beta}}_{\setminus j}^{t}$
Building on Algorithm 1, we outline the nodewise Multi-task Iterative Graphical Hard Thresholding (MIGHT) Algorithm 2 for the joint graphical estimation. For each node $j$, Algorithm 2 firstly constructs a column-scaled multi-task model (4), then applies Algorithm 1 to obtain a nodewise estimator, and finally recovers the joint model through the relationship (3). Notably, Algorithm 2 allows parallel computation across covariates (lines 6126-12), ensuring high computational efficiency.
Algorithm 2 Multi-task Iterative Graphical Hard Thresholding (MIGHT)
0:  $X^{(1)},\cdots,X^{(K)},~\kappa$
1:  for $k=1$ to K,K,~ do
2:   $\hat{\Sigma}^{(k)}=\frac{1}{n_{k}}\left(X^{(k)}\right)^{\top}X^{(k)}$
3:   $\hat{\Gamma}^{(k)}=\text{diag}\left(\{\hat{\Sigma}^{(k)}_{ii}\}_{i\in[p]}\right)$
4:   $Z^{(k)}=\sqrt{\frac{C_{0}N}{n_{k}}}X^{(k)}\left(\hat{\Gamma}^{(k)}\right)^{-1/2}$ //Column scaling
5:  end for
6:  for $j=1$ to p,p,~ do
7:   $\hat{\bm{\beta}}_{\setminus j}=\mathtt{Algorithm1}(Z_{\setminus j},X_{j},\kappa,\hat{s}_{0,j})$, with $\hat{s}_{0,j}$ tuned by (5) // Multi-task model
8:   for $k=1$ to K,K,~ do
9:    $\hat{\Theta}_{j,j}^{(k)}=n_{k}\cdot\left\|X_{j}^{(k)}-Z_{\setminus j}^{(k)}\hat{\bm{\beta}}_{\setminus j}^{(k)}\right\|_{2}^{-2}$
10:    $\hat{\Theta}_{\setminus j,j}^{(k)}=-\sqrt{\frac{C_{0}N}{n_{k}}}\hat{\Theta}_{j,j}^{(k)}\left(\hat{\Gamma}_{\setminus j,\setminus j}^{(k)}\right)^{-1/2}\hat{\bm{\beta}}_{\setminus j}^{(k)}$
11:   end for
12:  end for
12:  $\hat{\Theta}^{(1)},\cdots,\hat{\Theta}^{(K)}$
2.3 Miscellaneous setups
To ensure the adaptability, we employ a variant of the Birgé-Massart criterion, as introduced by Zhang et al. (2024), to determine the neighbor-similarity degree $s_{0,j}$ in Algorithm 1. Concretely, for each $j\in[p]$, let $\hat{\bm{\beta}}_{\setminus j}(s_{0,j})$ denote the estimation obtained under parameter $s_{0,j}$ in Algorithm 1. We define $\hat{s}(s_{0,j}):=\max\left\{1,~\left|\bigcup_{k\in[K]}\text{supp}\left(\hat{\bm{\beta}}_{\setminus j}^{(k)}(s_{0,j})\right)\right|\right\}$ as the estimated number of neighbors to node $j$, and $\hat{A}(s_{0,j}):=\left\|\hat{\bm{\beta}}_{\setminus j}(s_{0,j})\right\|_{0}$ as the total number of selected edges (to node $j$) among all $K$ graphs. We select $\hat{s}_{0,j}$ by the following information criterion:
$\hat{s}_{0,j}=$ $\arg\min_{s_{0,j}}\left\{\log\left(\frac{1}{N}\left\|X_{j}-Z_{\setminus j}\hat{\bm{\beta}}_{\setminus j}(s_{0,j})\right\|_{2}^{2}\right)\right.+\left.\frac{C}{N}\left[\hat{s}(s_{0,j})\cdot\log p+\hat{A}(s_{0,j})\cdot\log\left(K\cdot\hat{s}(s_{0})\right)\right]\right\}.$ (5)
Considering the symmetry of the precision matrix, we adopt the minimum symmetrization approach Cai et al. (2011, 2016); Shu et al. (2024):
$\hat{\Theta}_{i,j}^{(k),\text{Sym}}=\hat{\Theta}_{i,j}^{(k)}\cdot\mathbf{1}\left(\left|\hat{\Theta}_{i,j}^{(k)}\right|\leq\left|\hat{\Theta}_{j,i}^{(k)}\right|\right)+\hat{\Theta}_{j,i}^{(k)}\cdot\mathbf{1}\left(\left|\hat{\Theta}_{j,i}^{(k)}\right|<\left|\hat{\Theta}_{i,j}^{(k)}\right|\right),$
where $\hat{\Theta}^{(1)},\cdots,\hat{\Theta}^{(K)}$ are the estimated precision matrices obtained from Algorithm 2.
3 Theoretical analysis
This section provides the theoretical guarantee for our MIGHT method. For simplicity, we assume that the observations of all datasets are of the same order $n$, i.e., $n_{1}\asymp\cdots\asymp n_{k}\asymp n$, and $N=\sum_{k=1}^{K}n_{k}\asymp nK$. We make the following assumptions:
Assumption 1 (Gaussianity)
For each $k\in[K]$, we assume each row of $X^{(k)}$ independently drawn from $N(\mathbf{0}_{p},\Sigma^{(k)})$.
Assumption 2 (Bounded eigenvalues)
There exist two positive constants $\overline{C}>\underline{C}>0$ such that $\underline{C}\leq\min_{k\in[K]}\Lambda_{\min}\left(\Sigma^{(k)}\right)\leq\max_{k\in[K]}\Lambda_{\max}\left(\Sigma^{(k)}\right)\leq\overline{C}$.
Assumption 3 (Joint structure)
Assume the joint graphical model $\left(\Theta^{(1)},\cdots,\Theta^{(K)}\right)$ has the joint $(s_{j},s_{0,j})_{j\in[p]}$ structure, i.e., has both $\{s_{j}\}_{j\in[p]}$ local edge-sparsity and $\{s_{0,j}\}_{j\in[p]}$ local neighbor-similarity, with $1\leq s_{0,j}\leq K,~1\leq s_{j}\leq p-1$ for each $j\in[p]$.
Assumption 1 requires that the observations follow from Gaussian distributions, and this condition can be relaxed to sub-Gaussian cases. Assumption 2 excludes singular or nearly singular precision matrices and ensures that the variance in each nodewise regression, i.e., $\text{Var}(\epsilon_{j}^{(k)})=1/\Theta_{j,j}^{(k)}$, neither diverges nor approaches zero. This assumption is common in studies on graphical models Ren et al. (2015); Ma and Michailidis (2016). Assumption 3 characterizes the structure of the joint graphs, as we introduced in Definition 2 and 1, which indicates $|S_{j}|=s_{j}$ and $\sum_{k\in[K]}|S_{j}^{(k)}|=s_{j}s_{0,j}$ for every $j\in[p]$.
3.1 Estimation error bound
The following theorem provides column-wise bounds on the estimation error. For ease of display, for every $j\in[p]$ and $k\in[K]$, we define $\hat{S}_{j}^{(k)}:=\text{supp}\left(\hat{\Theta}_{\setminus j,j}^{(k)}\right)$ as the estimated neighbor set of node $j$ in the $k$-th graph.
Theorem 1 (Convergence rate and joint structure)
Assume Assumptions 1-3 hold and $n\gtrsim\max_{j\in[p]}\left(s_{j}\log p+s_{j}s_{0,j}\log K\right)$. We apply the MIGHT Algorithm 2 with taking
$C_{1}^{\text{alg}}\geq\frac{4\sqrt{3\overline{C}}\left(\kappa/\delta+\sqrt{3}-1\right)}{\kappa/\delta-1},\quad C_{2}^{\text{alg}}\geq\sqrt{\frac{\overline{C}^{2}(\overline{C}^{2}+\underline{C}^{2})}{\underline{C}^{3}}},$
where we define $\delta:=\frac{\overline{C}^{2}}{\overline{C}^{2}+\underline{C}^{2}}\in(0.5,1)$. Then, with probability at least $1-C_{3}e^{-C_{4}\log(pK)}$, we simultaneously have the following properties for every $j\in[p]$:
  1. 1.
    Non-asymptotic estimation error bound:
    $\sum_{k=1}^{K}\frac{n_{k}}{N}\left\|\hat{\Theta}^{(k)}_{\cdot,j}-\Theta^{(k)}_{\cdot,j}\right\|_{2}^{2}\lesssim$ $\frac{K+s_{j}\log p+s_{j}s_{0,j}\log(s_{j}K)}{N}.$ (6)
  2. 2.
    Sparse neighbor selection and edge selection: | _k [K] ^S_j^(k) | = O(s_j), and _k[K]| ^S_j^(k) | = O(s_j s_0,j).
The joint estimation bound in (6) consists of three components: (i) estimating the $K$ non-zero diagonal entries $\hat{\Theta}_{j,j}^{(k)}$, with rate $K/N$; (ii) estimating $s_{j}$ neighbors from p1p-1 nodes, with rate $(s_{j}\log p)/N$; and (iii) estimating $s_{j}s_{0,j}$ edges from $s_{j}K$ potential edges, with rate $(s_{j}s_{0,j}\log(s_{j}K))/N$. This decomposition demonstrates that our MIGHT method can adapt to the local $(s_{j},s_{0,j})$ joint structure of each node $j$, yielding an adaptive error bound while ensuring the estimator maintains joint sparsity. Additionally, by combining the node-specific bound (6), we can obtain the joint estimation bound in Frobenius norm as
$\sum_{k=1}^{K}\frac{n_{k}}{N}\left\|\hat{\Theta}^{(k)}-\Theta^{(k)}\right\|_{F}\lesssim\sqrt{\sum_{j\in[p]}\frac{K+s_{j}\log p+s_{j}s_{0,j}\log(s_{j}K)}{N}}.$ (7)
This result offers a more refined error bound than the separation estimation rate $\sqrt{K\cdot\sum_{j\in[p]}\frac{s_{j}\log p}{N}}$. It is also sharper than the bound in existing joint estimation methods Cai et al. (2016); Wang et al. (2017). Furthermore, even when prior support structure information is known, the upper bound derived by Ma and Michailidis (2016), $\sqrt{\frac{\sum_{j\in[p]}s_{j}\log p}{N}}+\sqrt{\frac{K\sum_{j\in[p]}s_{j}}{N}}$, could be less tight than ours in (7) in the case when $\sum_{j\in[p]}s_{j}>\sum_{j\in[p]}s_{j}\frac{s_{0,j}\log(s_{j}K)}{K}$. This scenario can be readily satisfied when the inter-graph local similarity is relatively low for some of the nodes. This demonstrates the superior accuracy and adaptability of our method.
3.2 Support recovery and asymptotic distribution
We further analyze the performance of our MIGHT estimator in exact support recovery (i.e., neighbor and edge selection consistency) and discuss its asymptotic properties. We begin by presenting the minimum signal conditions required for these results.
Assumption 4 (Minimum signal conditions for node $j$)
For a node $j\in[p]$, there exists a sufficiently large constant $C_{1}>0$ such that
$\min_{k\in[K]}\min_{i\in S_{j}^{(k)}}\left|\Theta^{(k)}_{i,j}\right|^{2}\geq$ $\frac{C_{1}}{n_{k}}\left(\frac{\log p}{s_{0,j}}+\log(s_{j}K)\right),$ (8)
$\min_{i\in S_{j}}\sum_{k\in[K]}\frac{n_{k}}{N}\left(\Theta^{(k)}_{i,j}\right)^{2}\geq$ $C_{1}\left(\frac{\log p+s_{0,j}\log(s_{j}K)}{N}\right).$ (9)
For a node $j\in[p]$, this assumption requires: (i) each nonzero per-graph edge $\Theta^{(k)}_{i,j}$ to exceed the per-graph noise level (8), and (ii) each neighbor $i\in S_{j}$ to have sufficiently large aggregated squared strength across graphs (9). They play complementary roles in guaranteeing exact edge- and neighbor-level support recovery. These two signal conditions are not implied by one another, because $s_{0,j}$ captures only the average occurrence for the neighbors across graphs, not the exact number of graphs in which a specific iji-j edge appears. Based on this assumption, Theorem 2 establishes the local selection consistency.
Theorem 2 (Selection consistency and sharp error bound)
[…]
  1. 1.
    Selection consistency: (^S_j^(1), , ^S_j^(K) ) = (S_j^(1), , S_j^(K) ).
  2. 2.
    Sharper estimation error bound: _k [K] nkN ∥^Θ^(k)_, j - Θ^(k)_, j _2^2 K + sjs0,j+ logpN.
Theorem 2 shows that our method is inherently signal-adaptive: it adaptively achieves support recovery and sharper estimation rates for those nodes with strong signals, and maintains the results in Theorem 1 for those nodes with relatively weak signals. This property is a distinct advantage of our dynamic regularized method over the existing joint estimation methods Ma and Michailidis (2016); Cai et al. (2016). Furthermore, define $\hat{S}_{j+}^{(k)}:=\hat{S}_{j}^{(k)}\cup\{j\}=\text{supp}\left(\hat{\Theta}_{\cdot,j}^{(k)}\right)$. The following theorem establishes the asymptotic property of nodes with strong signals.
Theorem 3 (Asymptotic normality)
Assume that all conditions in Theorem 2 hold, and that $\sqrt{n}\succ\max_{j\in\mathcal{A}}\left(s_{j}\log p\right)$. Then, as $n\to\infty$, for every $(j,k)\in\mathcal{A}\times[K]$, we have
$\max_{i\in\hat{S}_{j+}^{(k)}}~\sup_{w\in\mathbb{R}}\left|\mathbf{P}\left(\frac{\sqrt{n_{k}}\left(\hat{\Theta}^{(k)}_{i,j}-\Theta^{(k)}_{i,j}\right)}{\sigma_{i,\hat{S}_{j}}^{(k)}}\leq w\right)-\Phi(w)\right|\to 0,$
where $\Phi(\cdot)$ is the cumulative distribution function of standard normal distribution and $\sigma_{i,\hat{S}_{j}}^{(k)}$ is defined as $\left(\sigma_{i,\hat{S}_{j}}^{(k)}\right)^{2}:=\operatorname{Var}\left(\left(\Sigma^{(k)}_{\hat{S}_{j+}^{(k)},\hat{S}_{j+}^{(k)}}\right)^{-1}_{i,\cdot}x_{\hat{S}_{j+}^{(k)}}^{(k)}\left(x_{\hat{S}_{j+}^{(k)}}^{(k)}\right)^{\top}\left(\Sigma^{(k)}_{\hat{S}_{j+}^{(k)},\hat{S}_{j+}^{(k)}}\right)^{-1}_{\cdot,j}\Big|\hat{S}_{j}^{(k)}\right).$
Theorem 3 demonstrates that our MIGHT method possesses asymptotic normality, thereby enabling statistical inference. This property further strengthens the signal adaptivity of the MIGHT estimator, demonstrating its superiority of our method. In application, we estimate $\left(\sigma^{(k)}_{i,\hat{S}_{j}}\right)^{2}$ by using
$\frac{1}{n_{k}}\sum_{\ell=1}^{n_{k}}\left\{\left(\hat{\Sigma}^{(k)}_{\hat{S}_{j+}^{(k)},\hat{S}_{j+}^{(k)}}\right)^{-1}_{i,\cdot}X_{\ell,\hat{S}_{j+}^{(k)}}^{(k)}\left(X_{\ell,\hat{S}_{j+}^{(k)}}^{(k)}\right)^{\top}\left(\hat{\Sigma}^{(k)}_{\hat{S}_{j+}^{(k)},\hat{S}_{j+}^{(k)}}\right)^{-1}_{\cdot,j}\right\}^{2}-\left(\hat{\Theta}_{i,j}^{(k)}\right)^{2},$
where we denote by $X_{\ell,\hat{S}_{j+}^{(k)}}^{(k)}\in\mathbb{R}^{|\hat{S}_{j+}^{(k)}|}$ the $\ell$-th observation of the covariates indexed by $\hat{S}_{j+}^{(k)}$ in the $k$-th dataset.
4 Numerical experiments
This section evaluates the performance of our proposed method, MIGHT, against five existing methods: three joint estimation methods: the joint estimator method by Guo et al. (2011) (JEM), the group graphical lasso by Danaher et al. (2014) (GGL), and the fast and scalable joint estimator by Wang et al. (2017) (FJEM); and two separate estimation methods: the separate graphical lasso by Friedman et al. (2008) (Sep Glasso) and the separate nodewise $L_{0}$-regression by Shu et al. (2024) (Sep Node). We use the Bayesian Information Criterion (BIC) to select the optimal hyperparameters for these five methods. All code for this section has been uploaded to https://github.com/ShixiangLIUstat/MIGHT-Algorithm.
We set $p=100,K=10$ and $n_{1}=\cdots=n_{K}=100$. Data generation begins with an Erdős–Rćenyi base graph $G$, where each pair of nodes connects independently with a probability of $0.1$. We generate $K$ graphs $G_{1},\cdots,G_{K}$ by independently pruning a fraction $\rho$ of edges from the base graph $G$, with $\rho\in(0,1)$ controlling neighbor-similarity and edge-sparsity. The value of each edge in $G_{k}$ is uniformly sampled from $1,-0.5]\cup[0.5,1]$, forming the adjacency matrix $\Omega^{(k)}$. Then, the positive-definite precision matrix $\Theta^{(k)}$ is constructed as $\Theta^{(k)}=\Omega^{(k)}+\left(r+|\Lambda_{\min}(\Omega^{(k)})|\right)\cdot\mathbf{I}_{p}$, where parameter $r>0$ modulates the signal strength by inflating diagonal entries, thereby reducing the signal-to-noise ratio (see (3)) in the nodewise regression. Finally, by applying the corresponding covariance matrices $\left\{\Sigma^{(k)}\right\}_{k\in[K]}=\left\{(\Theta^{(k)})^{-1}\right\}_{k\in[K]}$, we get $K$ datasets as $X^{(1)},\cdots,X^{(K)}$.
To assess performance, we use the estimation error in Frobenius norm, defined as $\sum_{k=1}^{K}\frac{n_{k}}{N}\left\|\hat{\Theta}^{(k)}-\Theta^{(k)}\right\|_{F}$, and the column-wise error in the maximum squared $L_{2}$-norm (Max L2 Norm), defined as $\max_{j\in[p]}\sum_{k\in[K]}\frac{n_{k}}{N}\left\|\hat{\Theta}^{(k)}_{\cdot j}-\Theta^{(k)}_{\cdot j}\right\|_{2}^{2}$. Additionally, we use Matthews Correlation Coefficient (MCC) to evaluate the selection accuracy of edges in $K$ graphs (MCC-Edge), and of neighbor set $S_{j}=\bigcup_{k\in[K]}S_{j}^{(k)}$ for each node $j\in[p]$ (MCC-Ngbr).
4.1 Simulation 1: similarity and sparsity
We fix $r=0.1$ and vary $\rho\in\{0.2,0.5,0.8\}$, with each setting 100 times simulation. Table 1 presents the simulation results; all MCC values have been multiplied by 100, and standard errors are shown in parentheses. Our MIGHT method demonstrates superior performance across all evaluation metrics compared to competing methods. As $\rho$ increases, the recurrence frequency of shared neighbors across the $K$ graphical models decreases, leading to diminished structural similarity and a decline in the MCC-Ngbr metric. Additional performance metrics (e.g., Hamming Loss and F1 score) are provided in Appendix F in the supplementary material.
Table 1: The estimation performance across varying $\rho$.
Estimation Accuracy Support Recovery
Methods Frobenius Norm Max L2 Norm MCC-Edge MCC-Ngbr
$\rho=0.2$
MIGHT 18.979 (0.113) 6.257 (0.088) 63.662 (0.305) 89.673 (0.267)
GGL 23.428 (0.111) 8.772 (0.094) 45.276 (0.313) 39.520 (0.665)
JEM 25.272 (0.145) 9.854 (0.110) 54.262 (0.599) 67.863 (0.531)
FJEM 48.998 (0.116) 28.037 (0.152) 0.050 (0.050) 0.140 (0.140)
Sep Glasso 33.667 (0.151) 16.061 (0.127) 29.715 (0.303) 42.277 (0.279)
Sep Node 24.106 (0.104) 10.877 (0.135) 22.251 (0.198) 56.403 (0.459)
$\rho=0.5$
MIGHT 13.510 (0.127) 3.230 (0.075) 71.874 (0.488) 87.528 (0.432)
GGL 17.704 (0.183) 5.079 (0.110) 56.445 (0.760) 53.118 (1.872)
JEM 18.609 (0.154) 5.487 (0.099) 61.490 (0.767) 72.527 (0.757)
FJEM 39.491 (0.154) 18.167 (0.168) 0.099 (0.099) 0.183 (0.183)
Sep Glasso 25.434 (0.144) 9.410 (0.123) 40.385 (0.362) 47.693 (0.366)
Sep Node 18.046 (0.108) 6.092 (0.126) 31.874 (0.341) 62.222 (0.607)
$\rho=0.8$
MIGHT 7.290 (0.062) 1.056 (0.027) 81.258 (0.310) 83.614 (0.297)
GGL 9.610 (0.072) 1.577 (0.028) 55.902 (0.209) 58.346 (0.237)
JEM 10.210 (0.080) 1.811 (0.038) 69.879 (0.504) 74.619 (0.562)
FJEM 38.491 (4.866) 2727.8 (1556.5) 0.940 (0.236) 1.257 (0.369)
Sep Glasso 14.234 (0.076) 3.157 (0.045) 55.920 (0.269) 51.550 (0.202)
Sep Node 9.824 (0.066) 1.801 (0.036) 54.403 (0.381) 68.881 (0.430)
4.2 Simulation 2: signal strength
We fix $\rho=0.5$ and vary $r\in\{0.2,0.25,0.33,0.5,1\}$, with $1/r\in\{1,2,3,4,5\}$ measuring the signal strength. The results are presented in Figure 2. As the signal strength increases, the MIGHT method demonstrates lower estimation error and higher selection accuracy, confirming that enhanced signal strength contributes to better performance, consistent with our theoretical results.
Figure 2: Performance metrics with increasing signal strength ($1/r$), with each point $50$ times repetitions. All MCC values have been multiplied by 100. FJEM exhibits unstable estimation, and we exclude its excessively large estimation errors.
4.3 Simulation 3: asymptotic distribution
We fix $p=50$, $\rho=0.5$, $r=0.1$, and compare the asymptotic performance of MIGHT, GGL, JEM, and Sep Node methods across six specific matrix entries: three diagonal elements (Figure 3) and three off-diagonal elements (Figure 4). The results demonstrate that the MIGHT method achieves better asymptotic normality compared to the separate $\ell_{0}$ estimator (Sep Node), while GGL and JEM, two joint estimation methods based on M-estimator, fail to exhibit this property.
Figure 3: Histogram of (diagonal) $\sqrt{n}\left(\hat{\Theta}_{ii}^{(k)}-\Theta_{ii}^{(k)}\right)/\hat{\sigma}_{ii}^{(k)}$ among different methods with $300$ times repetitions. The orange curve represents the density of the standard normal distribution, while the black and orange vertical lines indicate the empirical mean and the population mean (0), respectively.
Figure 4: Histogram of (off-diagonal) $\sqrt{n}\left(\hat{\Theta}_{ij}^{(k)}-\Theta_{ij}^{(k)}\right)/\hat{\sigma}_{ij}^{(k)}$ among different methods with $300$ times repetitions.
5 Real data analysis
In this section, we apply our MIGHT method to a gene-expression cancer RNA-seq dataset, which is accessible via the UCI Machine Learning Repository Fiorini (2016). This dataset encompasses 801 patients and comprises 20,531 genes. The patients are categorized into five distinct types based on various tumor types, namely BReast CAncer (BRCA), KIdney Renal cell Carcinoma (KIRC), COlon ADenocarcinoma (COAD), LUng ADenocarcinoma (LUAD), and PRostate ADenocarcinoma (PRAD). The numbers of patients in these subtypes are 300, 78, 146, 141, and 136, respectively. Although these subtypes exhibit biological differences, they all involve tumor tissue, which necessitates the joint graphical model estimation.
5.1 The joint graphical model estimation
Our primary objective is to generate joint graphical representations of the relationships among the genes within each type, achieved by estimating their precision matrices. Out of the 20,531 genes, we perform a Kruskal–Wallis test on every gene and select the top 500 genes with the lowest p-values. Subsequently, our MIGHT method is used to jointly estimate the precision matrices for these five subtypes, with results shown in Figure 5. The black lines are the edges shared by some of the types, and the colored lines are the edges shared by only one specific type. Notably, most edges appear as black lines, indicating the common structures shared across multiple types of tumors.
Figure 5: Gene network for five types of tumors.
5.2 The classification tasks based on QDA
The estimation accuracy of our MIGHT method is then evaluated by applying a downstream classification task based on quadratic discriminant analysis (QDA) inspired by Price et al. (2015). We employ stratified sampling: each time, 80% of the subjects from each class are randomly selected to form the training set (640 samples), while the remaining samples constitute the testing set (161 samples). The quadratic discriminant scores are calculated as follows:
$\delta_{k}(\mathbf{x})=\log\hat{\pi}_{k}+\frac{1}{2}\log\left|\hat{\Theta}^{(k)}\right|-\frac{1}{2}(\mathbf{x}-\hat{{\mu}}^{(k)})^{\top}\hat{\Theta}^{(k)}(\mathbf{x}-\hat{{\mu}}^{(k)}),$ (10)
where $k=1,\cdots,5$, $\hat{\pi}_{k}=n_{k}/N$ is the proportion of observations belonging to the class $k$, and ${\hat{\mu}}_{k}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\bm{x}_{i}^{(k)}$ is the empirical mean for the training set of each class $k$. The classification rule is then given by $\hat{k}(\mathbf{x})=\arg\max_{k=1,\cdots,5}\delta_{k}(\mathbf{x})$.
The six estimation methods in Section 4 are used to compute quadratic discriminant scores (10) and perform classification predictions, respectively. Prediction performance is evaluated through the average TPR, FPR, Accuracy, and MCC. Table 2 summarizes the classification results. The classification based on the MIGHT method yields the best performance, demonstrating the superiority of our approach in this practical application.
Table 2: The classification performance for different methods with 50 repetitions.
Method TPR FPR Accuracy MCC
MIGHT 0.766 (0.037) 0.060 (0.009) 0.783 (0.034) 0.730 (0.043)
GGL 0.672 (0.027) 0.123 (0.008) 0.710 (0.024) 0.643 (0.032)
JEM 0.743 (0.021) 0.040 (0.005) 0.763 (0.018) 0.712 (0.022)
FJEM 0.266 (0.025) 0.179 (0.007) 0.336 (0.062) 0.305 (0.039)
Sep Glasso 0.513 (0.021) 0.112 (0.007) 0.619 (0.022) 0.491 (0.024)
Sep Node 0.542 (0.041) 0.102 (0.010) 0.648 (0.041) 0.531 (0.051)
6 Conclusion and discussion
This paper proposes an adaptive and computationally efficient method for the joint estimation of multiple graphical models. Theoretical analysis demonstrates that the proposed method achieves superior estimation accuracy and establishes selection consistency and asymptotic property. Both simulation studies and empirical analyses validate the superiority of our method.
Moreover, our joint-learning framework can be well-suited to modern applications. For example, in the cross-domain recommendation task, a user’s interactions with the same items across different domains are often correlated. Joint estimation can effectively identify these similar structures across domains, enabling cross-domain knowledge fusion through multi-graph neural networks Ouyang et al. (2021). Moreover, our framework shows broader applicability in multiple directed acyclic graphs, multimodal data analysis, and so on. We leave them for future exploration.
[…]
Appendix A Additional simulation results
This section presents supplementary results for the simulation studies in Section 4.1. All parameters are the same as in Section 4.1. In terms of estimation accuracy, we additionally evaluate the maximum column-wise $L_{1}$-norm (Max L1 Norm), defined as $\max_{j\in[p]}\sum_{k\in[K]}\frac{n_{k}}{N}\left\|\hat{\Theta}^{(k)}_{\cdot j}-\Theta^{(k)}_{\cdot j}\right\|_{1}$. Regarding support set recovery, we incorporate Precision, Recall, F1 score, and Hamming Loss, calculated as follows:
$\text{Precision}=$ $\frac{TP}{TP+FP},$
$\text{Recall}=$ $\frac{TP}{TP+FN},$
$\text{F1 score}=$ $\frac{2}{(\text{Precision})^{-1}+(\text{Recall})^{-1}},$
$\text{Hamming loss}=$ $FP+FN,$
where TP denotes true positives, FP denotes false positives, and FN denotes false negatives. We compute these four quantities at two levels: (i) edge-level selection accuracy across all $K$ graphs (reported as Precision-Edge, Recall-Edge, F1-Edge, Hamming-Edge), and (ii) neighbor-level selection accuracy using the union $S_{j}=\bigcup_{k\in[K]}S_{j}^{(k)}$ for each node $j\in[p]$ (reported as Precision-Ngbr, Recall-Ngbr, F1-Ngbr, Hamming-Ngbr). All reported values are averages over 100 times Monte Carlo replications. Higher Precision/Recall/F1 and lower Max L1 Norm/Hamming loss indicate better performance.
Table 3 presents the supplementary results. All Precision/Recall/F1 values have been multiplied by 100, and standard errors are shown in parentheses. Consistent with Table I in Section IV-A, across all examined values of $\rho$, our MIGHT method consistently attains superior estimation accuracy compared to competing methods. In terms of the edge-level and the neighbor-level support recovery, the MIGHT estimator preserves relatively high Precision, and attains the best overall performance on composite metrics such as F1 score and Hamming loss.
Table 3: Additional measurements across varying $\rho$.
Estimation Edge-wise Support Recovery Neighbor-wise Support Recovery
Methods Max L1 Norm Precision-Edge Recall-Edge F1-Edge Hamming-Edge Precision-Ngbr Recall-Ngbr F1-Ngbr Hamming-Ngbr
$\rho=0.2$
MIGHT 8.023 (0.112) 98.266 (0.068) 44.479 (0.443) 61.202 (0.424) 4412.000 (68.402) 98.797 (0.096) 83.739 (0.547) 90.619 (0.327) 170.067 (6.506)
GGL 10.588 (0.124) 49.007 (0.542) 53.213 (0.496) 50.926 (0.290) 8029.333 (112.060) 25.086 (1.176) 97.219 (0.980) 39.294 (1.058) 3040.933 (109.350)
JEM 9.527 (0.147) 94.766 (0.146) 36.176 (0.656) 52.268 (0.703) 5155.933 (83.513) 99.855 (0.040) 52.077 (0.800) 68.347 (0.700) 470.800 (10.502)
FJEM 310.986 (12.031) 8.130 (0.069) 95.529 (0.241) 14.983 (0.117) 84806.800 (299.661) 9.888 (0.086) 100.000 (0.000) 17.993 (0.143) 8921.067 (8.536)
Sep Glasso 13.169 (0.130) 48.674 (0.304) 24.573 (0.335) 32.618 (0.314) 7932.200 (75.785) 33.437 (0.311) 78.050 (0.533) 46.768 (0.283) 1740.800 (23.391)
Sep Node 11.835 (0.138) 96.191 (0.262) 5.954 (0.089) 11.210 (0.159) 7377.600 (69.475) 95.383 (0.305) 37.898 (0.459) 54.203 (0.492) 626.533 (9.288)
$\rho=0.5$
MIGHT 4.645 (0.089) 96.415 (0.088) 54.565 (0.634) 69.628 (0.521) 2341.600 (44.850) 98.776 (0.099) 79.159 (0.588) 87.850 (0.357) 215.800 (6.790)
GGL 6.917 (0.123) 52.231 (0.653) 67.697 (0.917) 58.802 (0.523) 4675.267 (95.499) 45.193 (1.611) 87.357 (1.218) 58.645 (1.412) 1301.933 (112.747)
JEM 6.142 (0.096) 88.103 (0.230) 42.651 (0.991) 57.292 (0.903) 3112.267 (60.960) 99.752 (0.074) 53.247 (0.943) 69.293 (0.805) 462.733 (11.435)
FJEM 305.408 (16.413) 5.180 (0.042) 96.528 (0.231) 9.831 (0.075) 87189.600 (392.454) 9.949 (0.078) 100.000 (0.000) 18.095 (0.128) 8915.067 (7.679)
Sep Glasso 8.716 (0.119) 49.300 (0.323) 37.201 (0.622) 42.309 (0.404) 4982.133 (51.580) 35.931 (0.345) 83.046 (0.634) 50.094 (0.311) 1633.600 (27.844)
Sep Node 7.340 (0.139) 97.050 (0.225) 10.790 (0.217) 19.401 (0.351) 4409.733 (41.578) 96.462 (0.278) 42.081 (0.688) 58.508 (0.672) 586.533 (10.120)
$\rho=0.8$
MIGHT 1.767 (0.032) 94.462 (0.184) 70.161 (0.496) 80.488 (0.335) 673.800 (12.248) 97.949 (0.148) 73.731 (0.460) 84.106 (0.294) 245.933 (4.865)
GGL 2.786 (0.041) 37.528 (0.486) 87.525 (0.429) 52.441 (0.432) 3163.067 (68.096) 45.296 (0.398) 87.855 (0.493) 59.707 (0.298) 1048.133 (14.790)
JEM 2.582 (0.037) 83.763 (0.404) 60.347 (0.774) 70.029 (0.449) 1020.933 (14.343) 99.659 (0.069) 59.727 (0.772) 74.607 (0.594) 357.533 (7.459)
FJEM 513.932 (36.266) 2.084 (0.019) 98.942 (0.172) 4.082 (0.037) 92248.667 (469.228) 8.918 (0.062) 100.000 (0.000) 16.374 (0.105) 9017.133 (6.150)
Sep Glasso 3.708 (0.044) 46.726 (0.306) 68.719 (0.602) 55.566 (0.243) 2178.333 (25.182) 37.463 (0.284) 87.771 (0.464) 52.471 (0.243) 1405.933 (19.041)
Sep Node 2.581 (0.039) 94.732 (0.312) 31.719 (0.382) 47.491 (0.442) 1388.600 (13.522) 93.738 (0.380) 53.858 (0.575) 68.365 (0.511) 439.267 (6.891)